# Large Multi-modal Models Can Interpret Features in Large Multi-modal Models

### v1

@@ 1 Introduction
In addition to methodological contributions, our case studies also offer unique insights into LMMs. Firstly, we identify emotional features in LMMs and demonstrate that these features enable LMMs to generate or share emotions. Previous works have highlighted the exceptional capabilities of LMMs in EQ assessments [39] and their ability to read emotions. We extend this investigation by exploring the emotions of LMMs and steering the model to express its own feelings. Secondly, we identify the causes of certain model behaviors and analyze potential reasons for undesired outcomes, such as hallucinations. Adjusting the relevant features can rectify the mistake. Thirdly, some features in LMMs align with those found in the human brain, suggesting that interpreting LMMs could be informative for understanding how the human brain integrates multisensory inputs. For example, we find features that are activated by an action and their corresponding OCR descriptions, which align with the invariant visual features found in the study by [32]. Specifically, our contributions can be summarized as follows:
  • We introduce the use of Sparse Autoencoders (SAEs) to analyze the open-semantic features of Large Multimodal Models (LMMs) and demonstrate, for the first time in multimodal domain, that the features learned by SAE in a smaller LMM can be effectively interpreted by a larger LMM.
  • We propose an automated pipeline for interpreting these features and further demonstrate that they can be used to steer model behaviors to address specific issues or induce desired outputs.
  • Additionally, we propose a method for identifying the underlying causes of model behaviors and offer an analysis of how to address these issues, highlighting the potential for developing more reliable LMMs.
@@ 2.2 Zero-shot Identification of Concepts
To quantify the relevance of a feature’s activation to a given concept, we first refine the descriptive text using language models to enhance conciseness. For instance, the verbose description ”The feature activates on the train tracks …” is condensed to ”Train tracks”. This refinement is performed with a smaller LLM to minimize computational expense and time. Following this, GroundingDino-SAM [22] is employed to generate a segmentation mask based on the succinct interpretation. Subsequently, we construct a composite mask incorporating every object detected in the image. The IoU score between the segmentation mask and our activation mask is then calculated, serving as the reference score for the feature’s relevance. An example of the evaluation process is demonstrated in Fig. 3 where the refined explanation is being sent to the GroundingDino-SAM [22] to produce the segmentation ground truth and calculate the IOU with the activated region.
@@ 3.2 Interpretaion Pipeline Evaluation
scene IOU ()($\uparrow$)( ↑ ) $0.007\pm 1\times 10^{-3}$ 0.20
CS ()($\uparrow$)( ↑ ) $18.1\pm 6\times 10^{-2}$ 24.4
object IOU ()($\uparrow$)( ↑ ) $0.005\pm 5\times 10^{-4}$ 0.19
CS ()($\uparrow$)( ↑ ) $18.2\pm 2\times 10^{-2}$ 24.0
part IOU ()($\uparrow$)( ↑ ) $0.007\pm 2\times 10^{-3}$ 0.21
CS ()($\uparrow$)( ↑ ) $18.1\pm 5\times 10^{-2}$ 23.5
material IOU ()($\uparrow$)( ↑ ) $0.01\pm 8\times 10^{-3}$ 0.39
CS ()($\uparrow$)( ↑ ) $18.1\pm 1\times 10^{-1}$ 24.1
texture IOU ()($\uparrow$)( ↑ ) $0.007\pm 2\times 10^{-3}$ 0.21
CS ()($\uparrow$)( ↑ ) $18.4\pm 6\times 10^{-2}$ 20.9
colour IOU ()($\uparrow$)( ↑ ) $0.005\pm 2\times 10^{-3}$ 0.10
CS ()($\uparrow$)( ↑ ) $19.6\pm 7\times 10^{-2}$ 20.3
Total IOU ()($\uparrow$)( ↑ ) $0.005\pm 2\times 10^{-4}$ 0.20
CS ()($\uparrow$)( ↑ ) $18.2\pm 1\times 10^{-2}$ 23.6
@@
Table 2: Consistency scores for our explanation, computed based on 100 samples per concept for GPT-4o consistency and 10 samples for human consistency, highlighting the agreement between GPT-4o-generated and human-generated explanations.
@@ Results
We present the consistency scores for each concept in Tab. 2. To evaluate the consistency of our explanations with the activated image regions, we employ GPT-4 as a judge and conduct a human study. For GPT consistency, we sample 100 test cases per concept, provided that the total number of samples for the concept exceeds 100. For human consistency, we manually label the correctness of 10 samples for each concept.
@@ 4 Probing into the Features
In previous section, we demonstrate that we are able to locate and interpret the visual features in a LMM using an automatic pipeline. However, what differs LMM from traditional vision model is its conversation, reasoning, and generalization ability across different modalities. On other words, we believe that the features inside LMM are open-semantic and should not be limited into the concepts in [4]. In this section, we probe into the features in our SAE and try to find out how different features contribute to the final result, and how can these features being used to steer model’s behavior in different scenarioes.
@@ 4.1 Case Studies of Emotion Feature
When interacting with humans, it is essential for the model to demonstrate empathy and the ability to understand human emotions. In [39], it has been shown that large multimodal models (LMMs) can perceive emotions and Emotional Quotient (EQ), enabling them to understand and resonate with human feelings. Building on this, we specifically investigate the relevant features, exploring how the model comprehends these features and how they influence its reasoning processes. Through examples of various image features and their effects on model responses, we demonstrate that LMMs are capable of: 1) Connecting emotional concepts between text and visual features, such as actions and behaviors; 2) Engaging with human emotions by adjusting the corresponding features to intervene in the reasoning process manually. 3) Response to concept that with invariant features between modalities.
Figure 5: The feature that relates to sad. We probe and find out the feature that activated on ”sad”. By clamping this feature, we can enforce the model to share the feeling of sad
Sad
We present a feature that may be related to the feeling of ”sadness” and explore the potential for enabling the model to share emotional responses. After probing and confirming that the feature aligns with ”sadness,” we investigate whether manipulating this feature could influence the model’s reasoning to simulate emotional responses. To test this, we use a simple prompt, ”What is your feeling right now?” and ask the assistant. Without steering, the model responds in a neutral, standard AI assistant tone, showing no emotion. However, when we clamp the ”sad” feature to a high value, the model responds with ”sad” as shown in Fig. 5
@@
Beyond simply experiencing a feeling, it is also essential for the LMM to interact and share emotions when presented with a specific scenario. Toward this goal, we use the same method to probe the feature associated with ”happiness” and provide an image depicting a joyful scenario. When asked about its feelings without steering, the model responds that it does not experience emotions. However, similar to the ”sad” feature, when we clamp the ”happy” feature, the model responds with expressions of happiness, as shown in Fig. 6. This demonstrates that the model’s reasoning process can be effectively influenced.
@@
We discovered an intriguing feature that links the text-based emotional concepts of greedy and hungry to visual representations of the action eating and the word hungry. We notice that the feature activates in response to the word hungry in the image, suggesting that it connects not only to the action eat but can also extend to broader concepts. To test this, we clamp this feature and prompt the model with ”Tell me a story about Alice and Bob”; the generated response revolves around themes of greed as shown in Fig. 7. This demonstrates that the model can reason from the visual action eat to a broader concept encompassing greedy and hungry with a unified view.
Figure 7: A feature that relates to the concept ”eat”. We further investigate about the concept behind this feature and find out that model can reason from a visual action ”eat” into the concept ”concept” and ”greedy”
@@ 4.2 Low Level Perception Features
One important distinction we observed in our features, as compared to those found in Large Language Models in previous studies [35, 6, 13], is the presence of numerous low-level visual features. These features typically represent basic concepts related to common visual elements and often exhibit high activation across most images. For example, these features may be associated with color, shape, or fundamental visual patterns, such as mesh structures or repetitive designs. In the example shown in Fig. 6, the feature associated with ”happy” is only the 78th most activated feature, with many other low-level concepts present as well. We showcase some of these features in Sec. 12, highlighting a key difference between LMMs and LLMs.
@@
Figure 9: Feature that relates to the text ”Barcelona”. By clamping this feature to high value, we intervene the reasoning steps and hallucination in Fig. 8 disappears.
Figure 10: The high attribution area of different images and on the text. For images, we observe that features with high attribution mostly activate on positions that relate to key information about the question. For text, we observe that the ”Bolivia” token contributes the most to the answer ”yes”
Figure 11: Feature that relates to the text ”Los”. We validate our assumption by finding another feature that relates to text and mitigate the hallucination.
@@ 4.3 Localizing the Cause for Model Behaviors
To study the cause for this output, we set the answer token $v_{c}=yes$ and $v_{b}=no$ and apply the algorithm with calculate per token contribution. By employing this strategy, we can filter out features that causing the model to be more favored to answer ”yes” instead of answer ”no”. To be specifically in this case, we mainly focus on two points: 1) How does the model reason from the image and is the model reason from the correct path? 2) If the model is paying the correct attention in the image, which part of the text part is causing the hallucination. To answer these two questions, we sort features by their attribution effect on image and text separately and observe their common part with high attribution.
@@
In Fig. 10, we visualize image patches for common areas with high attribution among the features. To do this, we first filtered out the top 10 features with the highest attribution towards the final output ”yes” and visualized their attribution map. Examining these top features reveals that they primarily contribute to tokens associated with text elements, such as map legends, country names, and other key visual details. This observation demonstrates that the model is effectively focusing on relevant areas of the image and has the ability to accurately identifying where to extract necessary information. However, even with the correct visual perception ability, the model still fail to produce the final answer.
@@
To further investigate the source of the incorrect answer, we continue visualizing the attribution of text tokens in the question. As shown in Fig. 10, the token ”Bolivia” contributes most to features with high attribution toward the answer ”yes.” Additionally, tokens like ”to” and ”the,” along with concepts such as ”Amazon Basin,” also have a positive attribution to the hallucinated answer ”yes.” This partially explains why the model responds with ”yes” instead of ”no,” even after extracting useful information from the image. While reasoning from visual features, the model is also influenced by text, leading it to approach the question with its pretrained knowledge.
@@ 4.4 Application of Model Steering on Hallucination
To address this issue, we first consider which features can encourage the model to focus on text tokens in the image rather than the question text. We hypothesize that clamping activations for certain OCR features related to image text may prompt the model to prioritize image-based features over question text features. We identify two such features that reduce hallucinations. In Fig. 9, we locate a feature linked to the word ”Barcelona” and clamp it to a high value, leading the model to rely on image information rather than general geographical knowledge. To validate this, we identify a feature related to the word ”Los” on traffic signs and similarly clamp it in Fig. 11, prompting the model to conclude that Bolivia is absent in the image. This shows that the model can use image information but may sometimes follow incorrect reasoning. With minimal intervention, we successfully guide the model to prioritize image information over question text.
@@ 5 Conclusion
In summary, we present an in-depth examination on the internal structure of the LMM and introduce an automated pipeline for interpreting its open-semantic features. Furthermore, we propose methods for steering the model’s behavior through these features and for identifying the sources of errors within the LMM. By systematically analyzing both the structural and functional aspects of the model, we aim to offer valuable insights into its interpretability and reliability. We hope that these findings will not only contribute to the advancement of research in this area but also stimulate further exploration into this area.
@@ 7 Limitations
8 Detail about Prompt
We detail the prompts used in different stages of the automated pipeline in this section. The prompt for zero-shot identification of concepts is provided in Tab. 3. For this task, we utilize the LLaVA-NeXT-OV-72B model [17]. To refine labels and categorize explanations using large language models (LLMs), we use the prompts detailed in Tabs. 5 and 6. Specifically, LLaMA-3.1-Instruct-8B is used for label refinement, while LLaMA-3.1-Instruct-70B is employed for categorizing explanations. For high-throughput performance, the models are served using SGLang [41].
@@ 8 Detail about Prompt
9 More Steering Examples
In this section, we provide more steering examples that we discover during experiments. We perform a large scale steering on the 5000 size features subset we choose and then filtered some interesting examples here. In Fig. 12, the feature activates on money and when this feature is clamped to 50, the model output a story about saving funds and by a house. In Fig. 13, when a feature relates to a feature that relate to speech, the model output a story about a man who is a speech master. In Fig. 14, we found a feature that relate to unix/linux and its steering effect would output a story about Ubuntu. More interestingly, in Fig. 15, though the model response on a visual ”chair” object, when steering this feature, model would output a story relates to ”chairman” instead of a ”chair”. Another example is that in Fig. 16, when steering this feature related to ”key” or ”keychain”, the model output a story about developing some analytic software.
10 CLIP-Score and IOU details

### v2

@@ 1 Introduction
In addition to methodological contributions, our case studies also offer unique insights into LMMs. Firstly, we identify emotional features in LMMs and demonstrate that these features enable LMMs to generate or share emotions. Secondly, we discover the low-level perception neurons (e.g., shape, texture, and color), object neurons, scene neurons, and invariant visual neurons. Secondly, previous works have highlighted the exceptional capabilities of LMMs in EQ assessments [41] and their ability to read emotions. We extend this investigation by exploring the emotions of LMMs and steering the model to express its own feelings. Thirdly, we identify the causes of certain model behaviors and analyze potential reasons for undesired outcomes, such as hallucinations. Adjusting the relevant features can rectify the mistake.
  • For the first time in the multimodal domain, we propose a pipeline to automatically interpret the vast and open-semantic features in LMMs. SAEs are adopted to disentangle these features into mono-semantic neurons, and another LMM is used to interpret the neurons.
  • This pipeline additionally enables us to steer model behaviors to induce desired outputs, identify the underlying causes of model behaviors, and offer an analysis of how to address these issues.
  • Our case studies also provide unique insights into LMMs. We discover unique neurons in LMMs, localize the causes of model behaviors, and steer the model to eliminate hallucinations.
@@ 2.2 Zero-shot Identification of Concepts
To quantify the relevance of a feature’s activation to a given concept, we first refine the descriptive text using language models to enhance conciseness. For instance, the verbose description “The feature activates on the train tracks …” is condensed to “Train tracks”. This refinement is performed with a smaller LLM to minimize computational expense and time. Following this, GroundingDino-SAM [24] is employed to generate a segmentation mask based on the succinct interpretation. Subsequently, we construct a composite mask incorporating every object detected in the image. The IoU score between the segmentation mask and our activation mask is then calculated, serving as the reference score for the feature’s relevance. An example of the evaluation process is demonstrated in Fig. 3 where the refined explanation is being sent to the GroundingDino-SAM [24] to produce the segmentation ground truth and calculate the IOU with the activated region.
@@ 3.2 Interpretaion Pipeline Evaluation
scene IOU $(\uparrow)$ $0.007\pm 1\times 10^{-3}$ 0.20
CS $(\uparrow)$ $18.1\pm 6\times 10^{-2}$ 24.4
object IOU $(\uparrow)$ $0.005\pm 5\times 10^{-4}$ 0.19
CS $(\uparrow)$ $18.2\pm 2\times 10^{-2}$ 24.0
part IOU $(\uparrow)$ $0.007\pm 2\times 10^{-3}$ 0.21
CS $(\uparrow)$ $18.1\pm 5\times 10^{-2}$ 23.5
material IOU $(\uparrow)$ $0.01\pm 8\times 10^{-3}$ 0.39
CS $(\uparrow)$ $18.1\pm 1\times 10^{-1}$ 24.1
texture IOU $(\uparrow)$ $0.007\pm 2\times 10^{-3}$ 0.21
CS $(\uparrow)$ $18.4\pm 6\times 10^{-2}$ 20.9
colour IOU $(\uparrow)$ $0.005\pm 2\times 10^{-3}$ 0.10
CS $(\uparrow)$ $19.6\pm 7\times 10^{-2}$ 20.3
Total IOU $(\uparrow)$ $0.005\pm 2\times 10^{-4}$ 0.20
CS $(\uparrow)$ $18.2\pm 1\times 10^{-2}$ 23.6
@@
Table 2: Consistency scores for our explanation for GPT-4o and human, highlighting the agreement between GPT-4o-generated and human-generated explanations.
@@ Results
We present the consistency scores for each concept in Tab. 2. To evaluate the consistency of our explanations with the activated image regions, we employ GPT-4 as a judge and conduct a human study. We evaluate GPT consistency using a total of 600 test cases, with 100 test cases per concept. For human consistency, we manually label the correctness of 60 test samples, with each sample verified by two human experts, resulting in 120 evaluations.
3.3 Cross Layer Ablations
LLaVA (8th) LLaVA (25th) LLaVA (32th) Random
IOU 0.30 0.31 0.40 0.005
CS 22.82 24.92 26.55 18.2
Table 3: Ablation studies across different LLaVA layers
BLIP (25th) Random (BLIP)
CS 28.01 17.71
Table 4: Evaluation the effect on different model struction
Cross-Layer Ablations
To strengthen our experiments, we conduct additional analyses across different layers of LLaVA-NeXT-LLaMA3-8B [17]. Due to resource limitations, we train only three separate SAEs with $k=4096$, each applied to low-level, mid-level, and high-level transformer layers, and analyze 100 features within the model. Our results consistently show a significant improvement compared to random baselines, suggesting that these phenomena are universal across different transformer layers. This finding validates the results of [37, 6]. Additionally, we observe that as layer depth increases, both IOU and Clip-Score improve. This phenomenon confirms our observations and aligns with the perspectives of [30, 39].
3.4 Model Architecture Ablation
Model Architecture Ablation
To further validate our results and assess their universality, we test our approach on Instruct-BLIP-7B [9], a Q-Former-based model [19], using a middle-layer. The SAE settings remain consistent with those in Sec. 3.3. Since BLIP image tokens are limited to only 32 tokens, we evaluate only the Clip-Score on Instruct-BLIP-7B, comparing it to a random baseline. As shown in Tab. 4, our results demonstrate that our method generalizes across different models.
@@ 4 Probing into the Features
In the previous section, we demonstrate that we are able to locate and interpret the visual features in an LMM using an automatic pipeline. However, what distinguishes LMM from the traditional vision model is its ability to talk, reason, and generalize between different modalities. In other words, we believe that the features inside LMM are open-semantic and should not be limited to the concepts in [4]. In this section, we probe into the features in our SAE and try to find out how different features contribute to the final result, and how these features being used to steer model’s behavior in different scenarios.
@@ 4.1 Case Studies of Emotion Feature
When interacting with humans, it is essential that the model demonstrate empathy and the ability to understand human emotions. In [41], it has been shown that large multimodal models (LMMs) can perceive emotions and Emotional Quotient (EQ), enabling them to understand and resonate with human feelings. Building on this, we specifically investigate the relevant features, exploring how the model comprehends these features and how they influence its reasoning processes. Through examples of various image features and their effects on model responses, we demonstrate that LMMs are capable of: 1) Connecting emotional concepts between text and visual features, such as actions and behaviors; 2) Engaging with human emotions by adjusting the corresponding features to intervene in the reasoning process manually. 3) Response to the concept that with invariant features between modalities.
@@
Beyond simply experiencing a feeling, it is also essential for the LMM to interact and share emotions when presented with a specific scenario. Toward this goal, we use the same method to probe the feature associated with “happiness” and provide an image depicting a joyful scenario. When asked about its feelings without steering, the model responds that it does not experience emotions. However, similar to the “sad” feature, when we clamp the “happy” feature, the model responds with expressions of happiness, as shown in Fig. 5. This demonstrates that the model’s reasoning process can be effectively influenced.
@@
We discovered an intriguing feature that links the text-based emotional concepts of greedy and hungry to visual representations of the action eating and the word hungry. We notice that the feature activates in response to the word hungry in the image, suggesting that it connects not only to the action eat but can also extend to broader concepts. To test this, we clamp this feature and prompt the model with “Tell me a story about Alice and Bob”; the generated response revolves around themes of greed as shown in Fig. 6. This demonstrates that the model can reason from the visual action eat to a broader concept encompassing greedy and hungry with a unified view.
Quantitative Results for Steering
Currently, there are no methods that can quantitatively evaluate steering effects, even in LLM literature [37]. The difficulty lies in the fact that quantitative evaluations of steering effects heavily depend on the prompts, and the existence of certain neurons could inhibit these effects. We took an initial try to use LLMs for evaluations. The prompt is “What do you see in this image?” and the given image feature a pure white background. We selected 100 object neurons, and only 11 of them exhibit a difference before and after steering. We used GPT-4 to assess the relevance in 1-10 (10 is highest like MT-Bench) and compared it with a random selection, as shown in Appendix D.
Figure 6: A feature that relates to the concept “eat”. We further investigate about the concept behind this feature and find out that model can reason from a visual action “eat” into the concept “concept” and “greedy”
@@ 4.2 Low Level Perception Features
One key distinction in our features, compared to those in Large Language Models [37, 6, 14], is the presence of numerous low-level visual features. These features represent basic visual concepts, such as color, shape, and patterns, and often exhibit high activation across images. For instance, in Fig. 5, the feature for “happy” ranks only 78th, with many low-level concepts also present. We highlight these features in Appendix H, underscoring a key difference between LMMs and LLMs.
@@
Figure 8: Feature that relates to the text “Barcelona”. By clamping this feature to high value, we intervene the reasoning steps and hallucination in Fig. 7 disappears.
Figure 9: The high attribution area of different images and on the text. For images, we observe that features with high attribution mostly activate on positions that relate to key information about the question. For text, we observe that the “Bolivia” token contributes the most to the answer “yes”
Figure 10: Feature that relates to the text “Los”. We validate our assumption by finding another feature that relates to text and mitigate the hallucination.
@@ 4.3 Localizing the Cause for Model Behaviors
To study the cause for this output, we set the answer token $v_{c}=yes$ and $v_{b}=no$ algorithm to calculate per-token contributions. This allows us to filter out features that bias the model toward answering “yes” over “no.” Specifically, we focus on two points: 1) Whether the model reasons correctly from the image, and 2) If the model attends correctly to the image, which text components cause hallucination. To address these, we sort features by their attribution effects on image and text separately and identify common high-attribution features.
@@
In Fig. 9, we visualize image patches for common areas with high attribution among the features. To do this, we first filtered out the top 10 features with the highest attribution towards the final output “yes” and visualized their attribution map. Examining these top features reveals that they primarily contribute to tokens associated with text elements, such as map legends, country names, and other key visual details. This observation demonstrates that the model is effectively focusing on relevant areas of the image and has the ability to accurately identifying where to extract necessary information. However, even with the correct visual perception ability, the model still fail to produce the final answer.
@@
To further investigate the source of the incorrect answer, we continue visualizing the attribution of text tokens in the question. As shown in Fig. 9, the token “Bolivia” contributes most to features with high attribution toward the answer “yes.” Additionally, tokens like “to” and “the,” along with concepts such as “Amazon Basin,” also have a positive attribution to the hallucinated answer “yes.” This partially explains why the model responds with “yes” instead of “no,” even after extracting useful information from the image. While reasoning from visual features, the model is also influenced by text, leading it to approach the question with its pretrained knowledge.
@@ 4.4 Application of Model Steering on Hallucination
To address this, we identify features that encourage the model to focus on image text rather than question text. We hypothesize that clamping activations of certain OCR features can shift the model’s focus to image-based features. We find two such features that reduce hallucinations. In Fig. 8, clamping a feature linked to “Barcelona” prompts the model to rely on image information instead of general knowledge. Similarly, in Fig. 10, clamping a feature related to “Los” on traffic signs leads the model to conclude Bolivia is absent. This demonstrates that with minimal intervention, the model can prioritize image information while sometimes following incorrect reasoning.
@@ 5 Conclusion
In summary, we analyze the internal structure of the LMM and introduce an automated pipeline for interpreting its open-semantic features. We also propose methods to steer the model’s behavior and identify error sources. By examining both structural and functional aspects, we provide insights into its interpretability and reliability, aiming to advance research and encourage further exploration.
6 Acknowledgments
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012, MOE-T2EP20223-0002), and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ Appendix B Limitations
Appendix C Detail about Prompt
We detail the prompts used in different stages of the automated pipeline in this section. The prompt for zero-shot identification of concepts is provided in Tab. 5. For this task, we utilize the LLaVA-NeXT-OV-72B model [18]. To refine labels and categorize explanations using large language models (LLMs), we use the prompts detailed in Tabs. 7 and 8. Specifically, LLaMA-3.1-Instruct-8B is used for label refinement, while LLaMA-3.1-Instruct-70B is employed for categorizing explanations. For high-throughput performance, the models are served using SGLang [43].
@@ Appendix C Detail about Prompt
Appendix D Qualitative Steering Experiments
Steering Random
GPT-4o Score 6.36 2.02
Table 9: GPT4 Score Evaluation of Steering Effects
In Tab. 9, we present the results of the steering evaluation for selected cases. Due to the high cost of large-scale steering evaluations, qualitative results are largely absent in the literature, including  [37]. To address this, we take an initial exploratory step by using an LLM to assess steering examples, demonstrating a potential solution.
Appendix E More Steering Examples
Figure 16: The feature that relates to sad. We probe and find out the feature that activated on “sad”. By clamping this feature, we can enforce the model to share the feeling of sad
Sad
We present a feature that may be related to the feeling of “sadness” and explore the potential for enabling the model to share emotional responses. After probing and confirming that the feature aligns with “sadness,” we investigate whether manipulating this feature could influence the model’s reasoning to simulate emotional responses. To test this, we use a simple prompt, “What is your feeling right now?” and ask the assistant. Without steering, the model responds in a neutral, standard AI assistant tone, showing no emotion. However, when we clamp the “sad” feature to a high value, the model responds with “sad” as shown in Fig. 16
In this section, we provide more steering examples that we discover during experiments. We perform a large scale steering on the 5000 size features subset we choose and then filtered some interesting examples here. In Fig. 11, the feature activates on money and when this feature is clamped to 50, the model output a story about saving funds and by a house. In Fig. 12, when a feature relates to a feature that relate to speech, the model output a story about a man who is a speech master. In Fig. 13, we found a feature that relate to unix/linux and its steering effect would output a story about Ubuntu. More interestingly, in Fig. 14, though the model response on a visual “chair” object, when steering this feature, model would output a story relates to “chairman” instead of a “chair”. Another example is that in Fig. 15, when steering this feature related to “key” or “keychain”, the model output a story about developing some analytic software.
Appendix F CLIP-Score and IOU details